Search This Blog


Thursday 2nd July 2009 ~ The Heat Intensifies

Despite the very high temperatures today, the X1 service seemed once again to manage remarkably well. There were actually a fair few swaps, but I don't think any of these can be attributed to vehicles malfunctioning in the heat. The second day of the Royal Norfolk Show seemed to affect eastbound services only and as a result several Lowestoft arrivals were held over to do the turn following that to which they were rostered. K17 was the first turn affected and this came in with 37579 but left westbound at 10.25 with 37568. You may recall that last night 568 was despatched from Lynn to rescue a stranded eastbound service with a broken windscreen. Well the stricken vehicle turned out to be 20353 on K18 (at least that what makes sense) as 353 was back at Lynn this morning and 568 was at Lowestoft. Following this services from Lowestoft ran as follows :

K18 in with 37576 out with 37579
K19 in with 37573 out with 37576
K1 in with 20353 out with 37573
K2 in with 20104 out with 20353
So you get the picture ? After this normality resumed with 37572 sticking on K5 and 20104 becoming spare. Later though 37566 on L10 was kept at Lowestoft and 20104 headed back west on the 16.25 to Peterborough.
So much for the Royal Norfolk Show delays, there were other swaps too. K4 with 20352 was halted at Lynn at 09.54 with replacement brake linings required, so 20118 was procured and worked the rest of the turn. This left King's Lynn short to cover Hunstanton schools turns, so 20106 which had been merrily plying its trade on Y9 was removed at Lynn at 12.24 and - at least it would seem, though confirmation required - replaced by a repaired 20352.
Y6 was not seen today so again reports are welcome, but it would seem most likely to have been 37563. B10Ms out today were 20103/4/5/6/18. 20103 was K15 (extremely unusual for this turn to be anything other than a Lynn vehicle, but especially notable being a LFT B10M) and 20105 did K16 which, ironically, ends at Lowestoft. Gemini 37575 passed its MOT today and is now available for traffic again.
20121 has been out on a test run and may need a new ecu, it should be next for attention once 20123 has received a new injector pipe on which work is ongoing. Sam reports that 20115 is no longer at Lowestoft and is on loan to Ipswich, so its hectic spell of X1 activity may yet prove to be a flash in the pan.

No comments: